OpenStudy (anonymous):

find exact value: tan (pi/6 + pi4) Using sum and difference formula but how do I set this up?

OpenStudy (anonymous):

\[\tan \left( \frac{ \pi }{ 6 } + \frac{ \pi }{ 4 }\right)\]